OpenStudy (anonymous):

aFor a summer fundraiser, Rob is doing manual labor to help several families do home improvement projects. He is tracking the amount of time he spends each week with a spreadsheet. If his spreadsheet shows that during the past two weeks, he spent a total of 11,040 minutes working, how many days and hours has he worked? a. 6 days and 6 hours b. 6 days and 16 hours c. 7 days 16 hours d. 7 days 7 hours Help !

jimthompson5910 (jim_thompson5910):

1 day = 24 hrs 1 hr = 60 min 24*1 hr = 24*60 min 24 hrs = 1440 min 1 day = 1440 min Do you see where to go from here?

OpenStudy (anonymous):

still need more help :/

jimthompson5910 (jim_thompson5910):

Ok, here's another way to do it Divide 11,040 into 60 to convert 11,040 minutes to hours: 11040/60 = 184 So 11040 minutes = 184 hours -------------------------------------------------------- Now divide this into 24 to get 184/24 = 7 remainder 16 This means that 184 hours = 7 days, 16 hours

jimthompson5910 (jim_thompson5910):

So 11,040 minutes = 7 days 16 hours

OpenStudy (anonymous):

Thanks so much ! (:

jimthompson5910 (jim_thompson5910):

you're welcome